OpenStudy (anonymous):

use l'hospital's rule where appropriate to find the following limit lim−−>∞ (1+x)^1/x

OpenStudy (anonymous):

limit is 1.

OpenStudy (anonymous):

how? sry...i'm a lil confused on this one.

OpenStudy (anonymous):

Factor out x from top and bottom.. get (1/x + 1) / 1.... limit to infinity.. gets us closer to undefined in the numerator.. meaning x = 0 in the numerator.. we get (und+1)/1 = treat und as zero... (0+1)/1 = 1/1 = 1